OpenStudy (anonymous):

MEDAL!!!!!!!!! Functions f(x) and g(x) are shown below:

OpenStudy (anonymous):

f(x) = 3x2 + 12x + 16 g(x)

OpenStudy (anonymous):

Using complete sentences, explain how to find the minimum value for each function and determine which function has the smallest minimum y-value

OpenStudy (welshfella):

for f(x) value of x at minimum is -b/2a ( that the a and b in the function ax^2 + bx + c) Then plug this into the formula for f(x) to get minimum value of f(x)

OpenStudy (anonymous):

g "(x) = -16sin(4x - π) > 0 for x = 5π/8, etc. So therefore, g(5π/8) = 2sin(π/4) + 4 = 4 + √2 > f(-2) = 4

OpenStudy (welshfella):

For g(x) the minimum is at x = pi/2 so plug this into formula for g(x) to get the required minimum.

OpenStudy (anonymous):

so f(x) isf '(x) = 6x + 12 =0 => x = -2

OpenStudy (welshfella):

oh ok - you are using calculus now plug -2 into f(x)

OpenStudy (anonymous):

that equals 0

OpenStudy (welshfella):

No - try again

OpenStudy (welshfella):

its 4

OpenStudy (anonymous):

what about g(x)

OpenStudy (anonymous):

is that the smallest y value

OpenStudy (welshfella):

yes g(x) is smallest minimum value

OpenStudy (welshfella):

i get minimum value of g(x) = -2

OpenStudy (welshfella):

g(x) = 2 sin (x - pi) g'(x) = 2 cos ( x - pi) = 0 at turning points

OpenStudy (anonymous):

explain how to find the minimum value for each function

OpenStudy (welshfella):

using calculus you find the first derivative and equate it to zero and solve for x this will give you a critical point you can now test for maximum or minimum by taking the second derivative and finding its sign for the value of x you have found positive value is minimum and negative is a maximum.

OpenStudy (welshfella):

so for the first question f"(x) is 6 positive so thats a minimum

OpenStudy (anonymous):

thanks for the help man

OpenStudy (welshfella):

yw